### v1

# Center of distances of ultrametric spaces generated by labeled trees

Abstract. The center of distances of a metric space $(X,d)$ is the set $C(X)$ of all $t\in\mathbb{R}^{+}$ for which the equation $d(x,p)=t$ has a solution for each $p\in X$. We prove that the equalities $C(X)=\{0\}$ or $C(X)=\{\operatorname{diam}X,0\}$ hold if $(X,d)$ is an ultrametric space generated by labeled trees. The necessary and sufficient conditions under which $\operatorname{diam}X\in C(X)$ are found.
Keywords: center of distances; diametrical graph; labeled tree; ultrametric space.
MSC (2020): 54E35.
1 Introduction
Ultrametric spaces generated by non-negative vertex labelings on arbitrary trees were introduced in [19] and subsequently studied in [11, 13]. A characterization of totally bounded ultrametric spaces generated by labeled almost rays was provided in [18]. Furthermore, [16] contains a metric characterization of ultrametric spaces generated by labeled star graphs.
It was shown in [10] that the compact ultrametric spaces generated by labeled star graphs are the completions of totally bounded ultrametric space generated by decreasingly labeled rays. The labeled trees generating separable ultrametric spaces are characterized, up to isomorphism, in [15].
In the present paper we investigate the class $\bf UT$} of all ultrametric spaces which are isometric to spaces generated by labeled trees. The main goal of the paper is the characterization of the center of distances $C(X)$ of $\bf UT$}-spaces $(X,d)$.
The paper is organized as follows. The next section contains some definitions and facts from the theory of ultrametric spaces and graph theory. The main results of the paper are proved in Sections 35.
Theorem 3 of Section 3 shows that the center of distances $C(X)$ of arbitrary $\bf UT$}-space $(X,d)$ satisfies either the equality
$C(X)=\{0,\operatorname{diam}X\}$ (1)
or the equality
$C(X)=\{0\}.$ (2)
In Section 4 we introduce the concept of centered spheres and use it in Theorem 4 to describe the geometry of $\bf UT$}-spaces $(X,d)$ which satisfy (1). Theorem 5 of this section gives a necessary and sufficient condition under which each open ball of a given $\bf UT$}-space is a centered sphere in this space. Theorem 6 of Section 4 shows that an ultrametric space $(X,d)$ is equidistant if and only if each centered sphere of $(X,d)$ is an open ball of $(X,d)$.
Theorem 7 of Section 5 shows that equality (1) holds for $\bf UT$}-space $(X,d)$ if and only if the diametrical graph of $(X,d)$ has a spanning star subgraph. Corollary 4 of Theorem 8 of Section 5 shows that equality (2) holds for $\bf UT$}-space $(X,d)$ with $|X|\geq 2$ if and only if $(X,d)$ is weakly similar to an unbounded ultrametric space.
2 Preliminaries
In what follows we will use the symbol $\mathbb{R}^{+}$ to denote the set $0,+\infty)$.
Definition 1.
Let $X$ be a non-empty set. A metric on $X$ is a symmetric function $d\colon X\times X\to\mathbb{R}^{+}$ satisfying the equality $d(x,x)=0$ for each $x\in X$, the inequality $d(x,y)>0$ for all distinct $x,y\in X$, and the triangle inequality
$d(x,y)\leq d(x,z)+d(z,y)$
for all $x$, $y$, $z\in X$.
A metric space $(X,d)$ is called an ultrametric space if the strong triangle inequality
$d(x,y)\leq\max\{d(x,z),d(z,y)\}$
holds for all $x$, $y$, $z\in X$. The function $d$ is called an ultrametric on $X$.
Following [7, p. 46] we say that a metric $d:X\times X\to\mathbb{R}^{+}$ is equidistant if there is a constant $k>0$ such that
$d(x,y)=k$
whenever $x$ and $y$ are distinct points of $X$.
Definition 2.
Let $(X,d)$ and $(Y,\rho)$ be metric spaces. A bijective mapping $\Phi\colon X\to Y$ is said to be an isometry if
$d(x,y)=\rho(\Phi(x),\Phi(y))$
holds for all $x,y\in X$. The metric spaces are isometric if there is an isometry of these spaces.
Let $(X,d)$ be a metric space. In what follows we will denote by $D(X)$ the distance set of the space $(X,d)$,
$D(X):=\{d(x,y):x,y\in X\}.$ (3)
The concept of center of distances was introduced in [4] as follows.
Definition 3.
Let $(X,d)$ be a metric space. The center of distances of $(X,d)$ is the set $C(X)$ defined as
$C(X):=\{t\in D(X):\forall p\in X\ \exists x\in X\ d(p,x)=t\}.$ (4)
Remark 1.
Some properties of the center of distances are described in [2, 3, 24, 25].
Example 1 (The $p$-adic ultrametric).
Let $p\geq 2$ be a prime number, $\mathbb{Q}$ be the set of rational numbers, and $\mathbb{Z}$ be the set of integers. The $p$-adic valuation of $t\in\mathbb{Q}$ is defined as
$|t|_{p}:=\begin{cases}p^{-\gamma},&\text{if }t\neq 0,\\ 0,&{\text{o}therwise},\end{cases}$ (5)
where $\gamma=\gamma(t)$ is the unique $\gamma\in\mathbb{Z}$ such that
$t=p^{\gamma}\frac{m}{n},$
and $m,$ $n$ are coprime to $p$. The $p$-adic valuation satisfies the inequality
$|t+w|_{p}\leq\max\{|t|_{p},|w|_{p}\}$
for all $t,w\in\mathbb{Q}$. This implies that the mapping $d_{p}:\mathbb{Q}\times\mathbb{Q}\to\mathbb{R}^{+},$
$d_{p}(t,w):=|t-w|_{p}$ (6)
is an ultrametric on $\mathbb{Q}$. (See formula (1.10) in Chapter 1 of [1].) It directly follows from (6) that the mapping
${\mathbb{Q}}\ni x\mapsto x+y\in{\mathbb{Q}}$
is a self-isometry of the space $({\mathbb{Q}},d_{p})$ for every $y\in\mathbb{Q}$. Which gives us the equalities
$D(\mathbb{Q})=C(\mathbb{Q})=\{\,p^{\gamma}:\gamma\in\mathbb{Z}\,\}\cup\{0\}$ (7)
in accordance with formulas (3)–(6).
Let $A$ be a non-empty subset of a metric space $(X,d)$. The quantity
$\operatorname{diam}A:=\sup\{d(x,y):x,y\in A\}$ (8)
is the diameter of $A$. If the inequality $\operatorname{diam}A<\infty$ holds, then we say that $A$ is a bounded subset of $(X,d)$.
Let $(X,d)$ be a metric space. The open ball with a radius $r>0$ and a center $c\in X$ is the set
$B_{r}(c):=\{x\in X\colon d(c,x)<r\}.$ (9)
Let us denote by ${\bf B}_{X}$ the set of all open balls of metric space $(X,d)$.
The next proposition is a part of Proposition 18.4 from [27].
Proposition 1.
Let $(X,d)$ be an ultrametric space. Then, for every ball $B_{r}(c)\in{\bf B}_{X}$ and every $a\in B_{r}(c)$, we have
$B_{r}(c)=B_{r}(a).$
Now, we recall the concepts of compactness and total boundedness.
A standard definition of compactness is usually formulated as: a metric space $(X,d)$ is compact if every open cover of $X$ has a finite subcover.
Definition 4.
A metric space $(X,d)$ is totally bounded if for every $r>0$ there is a finite set $\{B_{r}(x_{1}),\ldots,B_{r}(x_{n})\}\subseteq{\bf B}_{X}$ such that
$X\subseteq\bigcup_{i=1}^{n}B_{r}(x_{i}).$
It follows directly from Definition 4 that every compact metric space is totally bounded.
The next proposition is a corollary of Theorem 3.6 from paper [17].
Proposition 2.
Let $(X,d)$ be an infinite totally bounded ultrametric space and let $D(X)$ be the distance set of the space $(X,d)$. Then there is a strictly decreasing sequence $(x_{n})_{n\in\mathbb{N}}$ of points of $\mathbb{R}^{+}$ such that
$\lim_{n\to\infty}x_{n}=0$
holds and the equivalence
$(x\in D(X))\iff(x=0\text{ or }\exists n\in\mathbb{N}:x_{n}=x)$
is valid for every $x\in\mathbb{R}^{+}$.
Further, let us recall some definitions and facts from graph theory.
A graph is a pair $(V,E)$ consisting of a non-empty set $V$ and a set $E$ whose elements are unordered pairs $\{u,v\}$ of different points $u,v\in V$. For a graph $G=(V,E)$, the sets $V=V(G)$ and $E=E(G)$ are called the set of vertices and the set of edges, respectively. A graph $G$ is finite if $V(G)$ is a finite set. If $\{x,y\}\in E(G)$, then the vertices $x$ and $y$ are ends of $\{x,y\}$ and called adjacent. In what follows, we will always assume that $E(G)\cap V(G)=\emptyset$. A graph $G$ is empty if $E(G)=\emptyset$.
The following concept is well-known for finite graphs (see, for example, [8, p. 17]).
Definition 5.
Let $G$ be a graph and let $k\geq 2$ be a cardinal number. The graph $G$ is complete $k$-partite if the vertex set $V(G)$ can be partitioned into $k$ non-empty disjoint parts, in such a way that no edge has both ends in the same part and any two vertices in different parts are adjacent.
We say that $G$ is a complete multipartite graph if $G$ is complete $k$-partite for some $k$.
Our next definition is a modification of Definition 2.1 from [26].
Definition 6.
Let $(X,d)$ be a metric space. Denote by $G_{X}$ a graph such that $V(G_{X})=X$ and, for $u,v\in V(G_{X})$,
$(\{u,v\}\in E(G_{X}))\iff(d(u,v)=\operatorname{diam}X\text{ and }u\neq v).$ (10)
We call $G_{X}$ the diametrical graph of $(X,d)$.
Remark 2.
It directly follows from (10) that $G_{X}$ is empty if and only if $X$ is a singleton or $\operatorname{diam}X\notin D(X)$.
The following theorem is a direct consequence of Theorems 3.1 and 3.2 from [9].
Theorem 1.
Let $(X,d)$ be an ultrametric space with $|X|\geq 2$. Then the following statements are equivalent:
  1. (i)
    The diametrical graph $G_{X}$ is non-empty.
  2. (ii)
    The diametrical graph $G_{X}$ is complete multipartite.
Corollary 1.
Let $(X,d)$ be an ultrametric space with $|X|\geq 2$, let $D(X)$ be the distance set of $(X,d)$, and let $C(X)$ be the center of distances of $(X,d)$. If $\operatorname{diam}X$ belongs to $D(X)$,
$\operatorname{diam}X\in D(X),$ (11)
then $\operatorname{diam}X$ also belongs to $C(X)$,
$\operatorname{diam}X\in C(X).$ (12)
Proof.
If (11) holds, then the diametrical graph $G_{X}$ is non-empty by Remark 2. Hence $G_{X}$ is complete multipartite by Theorem 1. Now Definitions 3 and 5 imply (12). ∎
Let $G$ be a graph. A graph $G_{1}$ is a subgraph of $G$ if
$V(G_{1})\subseteq V(G)\quad\text{and}\quad E(G_{1})\subseteq E(G).$
In this case we will write $G_{1}\subseteq G$.
A path is a finite graph $P$ whose vertices can be numbered without repetitions so that
$V(P)=\{x_{1},\ldots,x_{k}\}\quad\text{and}\quad E(P)=\{\{x_{1},x_{2}\},\ldots,\{x_{k-1},x_{k}\}\}$ (13)
with $k\geqslant 2$. We will write $P=(x_{1},\ldots,x_{k})$ or $P=P_{x_{1},x_{k}}$ if $P$ is a path satisfying (13) and said that $P$ is a path joining $x_{1}$ and $x_{k}$. A graph $G$ is connected if for every two distinct vertices of $G$ there is a path $P\subseteq G$ joining these vertices.
A finite graph $C$ is a cycle if there exists an enumeration of its vertices without repetitions such that $V(C)=\{x_{1},\ldots,x_{n}\}$ and
$E(C)=\{\{x_{1},x_{2}\},\ldots,\{x_{n-1},x_{n}\},\{x_{n},x_{1}\}\}\quad\text{with }n\geq 3.$
Definition 7.
A connected graph $T$ without cycles is called a tree.
Let us recall the concept of star graphs.
Definition 8.
A tree $T$ is a star graph if there exists a vertex $c\in V(T)$ such that
$E(T)=\{\{v,c\}\colon v\in V(T)\setminus\{c\}\}.$
A subtree $T$ of a connected graph $G$ is called a spanning tree of $G$ if the equality $V(T)=V(G)$ holds.
Let us introduce the concept of labeled trees.
Definition 9.
A labeled tree is a pair $(T,l)$, where $T$ is a tree and $l$ is a mapping defined on $V(T)$.
If $(T,l)$ is a labeled tree, then we write $T=T(l)$ instead of $(T,l)$. Moreover, in what follows, we will consider only the non-negative real-valued labelings $l\colon V(T)\to\mathbb{R}^{+}$.
Following [19], for arbitrary labeled tree $T=T(l)$, we define a mapping $d_{l}\colon V(T)\times V(T)\to\mathbb{R}^{+}$ as
$d_{l}(u,v):=\begin{cases}\max\limits_{w\in V(P_{u,v})}l(w),&\text{if}\,\,u\neq v,\\ 0,&\text{otherwise},\end{cases}$ (14)
where $P_{u,v}$ is a path joining $u$ and $v$ in $T$.
Theorem 2.
Let $T=T(l)$ be a labeled tree. The mapping $d_{l}$ is an ultrametric on the set $V(T)$ if and only if the inequality
$\max\{l(u),l(v)\}>0$ (15)
holds for every $\{u,v\}\in E(T)$.
A proof of Theorem 2 can be obtained by simple modification of the proof of Proposition 3.2 in [19].
As in [18, 16] we shall say that a labeling $l\colon V(T)\to\mathbb{R}^{+}$ is non-degenerate if inequality (15) holds for every $\{u,v\}\in E(T)$.
The following lemma will be used in the next section of the paper.
Lemma 1.
Let $T=T(l)$ be a labeled tree with a non-degenerate labeling $l\colon V(T)\to\mathbb{R}^{+}$ and let $D(V(T))$ be the distance set of the ultrametric space $(V(T),d_{l})$. Then there exists a non-degenerate labeling $l^{*}\colon V(T)\to\mathbb{R}^{+}$ such that
$d_{l^{*}}=d_{l}$ (16)
and the equality
$D(V(T))=\{l^{*}(u)\,:\,u\in V(T)\}$ (17)
holds.
Proof.
Let us define a labeling $l^{*}\colon V(T)\to\mathbb{R}^{+}$ as
$l^{*}(u):=\begin{cases}l(u),&\text{if}\,\,l(u)\in D(V(T)),\\ 0,&\text{otherwise}.\end{cases}$ (18)
Equalities (16) and (17) evidently hold if $|V(T)|=1$. Suppose that $|V(T)|\geq 2$ and consider an arbitrary edge $\{u^{\circ},v^{\circ}\}\in E(T)$. Using (14) we obtain
$0<d_{l}(u^{\circ},v^{\circ})=\max\{l(u^{\circ}),l(v^{\circ})\}$ (19)
and, consequently,
$\max\{l(u^{\circ}),l(v^{\circ})\}\in D(V(T)).$ (20)
Membership relation (20) and formula (18) imply the equality
$\max\{l(u^{\circ}),l(v^{\circ})\}=\max\{l^{*}(u^{\circ}),l^{*}(v^{\circ})\}.$ (21)
The last equality and formula (18) show that $l^{*}\colon V(T)\to\mathbb{R}^{+}$ is a non-degenerate labeling of $V(T)$ because $l\colon V(T)\to\mathbb{R}^{+}$ is non-degenarate. Thus $(V(T),d_{l^{*}})$ is an ultrametric space by Theorem 2.
Let us prove equality (16). Let $x$ and $y$ be two distinct vertices of $T$ and let $P_{x,y}\subseteq T$ be the path joining $x$ and $y$. Then using (14), (19), and (21) we obtain
$d_{l}(x,y)=\max_{w\in V(P_{x,y})}l(w)=\max_{\{u,v\}\in E(P_{x,y})}\{l(u),l(v)\}=\max_{\{u,v\}\in E(P_{x,y})}\{l^{*}(u),l^{*}(v)\}=$
$\max_{w\in V(P_{x,y})}\{l^{*}(w)\}=d_{l^{*}}(x,y).$
Thus equality (16) holds.
To prove equality (17) it suffices to note that equality (14) with $l=l^{*}$ implies the inclusion
$D(V(T))\subseteq\{l^{*}(u):u\in V(T)\}$
and that the converse inclusion
$D(V(T))\supseteq\{l^{*}(u):u\in V(T)\}$
follows from (18).
The proof is completed. ∎
In what follows we will denote by $\bf UT$} the class of all ultrametric spaces which are isometric to the space generated by labeled trees with non-degenerate vertex labelings.
The main goal of the present paper is to describe the structure of centers of distances $C(X)$ for $(X,d)\in{\bf UT}$.
3 Center of distances of UT-spaces
Let $(X,d)$ be a metric space, let $p\in X$, and let $D_{p}(X)$ be a subset of the distance set $D(X)$ defined as
$D_{p}(X):=\{d(p,x):x\in X\}.$ (22)
Proposition 3.
Let $(X,d)$ be a metric space and let $C(X)$ be the center of distances of $(X,d)$. Then the equality
$C(X)=\bigcap_{p\in X}D_{p}(X)$ (23)
holds.
Proof.
All metric spaces are non-empty by Definition 1. Consequently, the right side of equality (23) is correctly defined. Now (23) follows from Definition 3 and formulas (4), (22). ∎
Proposition 4.
Let $(X,d)$ be a metric space and let $C(X)$ be the center of distances of $(X,d)$. Then $C(X)$ contains the point 0.
Proof.
Since we have $d(p,p)=0$ for every $p\in X$, equality (22) implies
$0\in D_{p}(X)$
for all $p\in X$. Hence
$0\in\bigcap_{p\in X}D_{p}(X)$
holds. The last membership relation and Proposition 3 imply $0\in C(X)$ as required. ∎
Remark 3.
Propositions 3 and 4 are known, see, for example, [25] and [4].
Lemma 2.
Let $T=T(l)$ be a labeled tree with a non-degenerate labeling $l\colon V(T)\to\mathbb{R}^{+}$ satisfying the equality
$D(V(T))=\{l(u):u\in V(T)\}.$ (24)
Then
$\alpha\notin C(X)$ (25)
holds whenever
$0<\alpha<\operatorname{diam}V(T).$ (26)
Proof.
Let us consider an arbitrary $\alpha$ satisfying (26). Then, using (8), (24), and (26), we can find a vertex $v^{*}\in V(T)$ such that
$\alpha<l(v^{*}).$ (27)
Write
$D_{v^{*}}(V(T)):=\{d_{l}(v^{*},u):u\in V(T)\}.$
Equality (14) implies the inequality
$d_{l}(v^{*},u)\geq l(v^{*})$
for each $u\in V(T)\setminus\{v^{*}\}$. Hence $\alpha$ does not belong to the set $D_{v^{*}}(V(T))$ by (27). Consequently (25) holds by Proposition 3. ∎
The next theorem completely describes the center of distances for ultrametric spaces $(X,d)\in{\bf UT}$.
Theorem 3.
Let $(X,d)\in{\bf UT}$ have at least two points, let $D(X)$ be the distance set of $(X,d)$, and let $C(X)$ be the center of distances of $(X,d)$. Then we have either
$C(X)=\{0\},$ (28)
or
$C(X)=\{0,\operatorname{diam}X\}$ (29)
and, moreover, the last equality holds if and only if
$\operatorname{diam}X\in D(X).$ (30)
Proof.
Let us prove the equivalence (29) $\Leftrightarrow$ (30).
Since $C(X)$ is a subset of $D(X)$, equality (29) implies membership relation (30).
Suppose now that we have (30). Then Corollary 1 implies $\operatorname{diam}X\in C(X).$ Proposition 4 gives us $0\in C(X).$ Consequently, the inclusion
$C(X)\supseteq\{0,\operatorname{diam}X\}$ (31)
is valid. Since $(X,d)$ belongs to the class $\bf UT$}, there exists a labeled tree $T=T(l)$ such that $(X,d)$ and $(V(T),d_{l})$ are isometric. By Lemma 1 we can also suppose that
$D(V(T))=\{\,l(u)\colon u\in V(T)\,\}.$ (32)
Then using Lemma 2 we obtain $C(V(T))\subseteq\{0,\operatorname{diam}V(T)\}.$ Since $(X,d)$ and $(V(T),d_{l})$ are isometric, the last inclusion gives the inclusion
$C(X)\subseteq\{0,\operatorname{diam}X\}.$ (33)
Equality (29) follows from (31) and (33). Thus the equivalence (29) $\Leftrightarrow$ (30) is valid.
Let us now assume that equality (29) does not hold. Then, as was shown above, the formula
$\operatorname{diam}X\notin D(V(T))$ (34)
is valid. Let $T=T(l)$ be a labeled tree such that $(X,d)$ and $(V(T),d_{l})$ are isometric and equality (32) holds. Then using Proposition 4, Lemma 2 and (34) we see that
$C(V(T))=\{0\}.$ (35)
Since $(X,d)$ and $(V(T),d_{l})$ are isometric, (28) follows from (35).
The proof is completed. ∎
Remark 4.
If $(X,d)$ contains exactly one point, then we have $\operatorname{diam}X=0$ and, consequently, equalities (28) and (29) are equivalent.
Example 2 (The Delhommé-Laflamme-Pouzet-Sauer ultrametric).
Let us define a mapping $d^{+}:\mathbb{R}^{+}\times\mathbb{R}^{+}\to\mathbb{R}^{+}$ as
$d^{+}(p,q):=\begin{cases}\max\{p,q\},&\text{if }p\neq q,\\ 0,&\text{otherwise}.\end{cases}$
Then $d^{+}$ is an ultrametric on $\mathbb{R}^{+}$ and the inequality
$d^{+}(0,p)\leq d(p,q)$
holds whenever $0\neq p\neq q$. The last inequality and Theorem 2.1 from [16] imply the existence of a labeled star graph $S=S(l)$ such that $V(S)=\mathbb{R}^{+}$ and $d^{+}=d_{l}$, where $d_{l}$ is an ultrametric defined by (14) with $T=S$. Thus $(\mathbb{R}^{+},d^{+})$ is an UT-space. Since the ultrametric space $(\mathbb{R}^{+},d^{+})$ is unbounded, Theorem 3 implies that the center of distances of $(\mathbb{R}^{+},d^{+})$ contains the unique point 0,
$C(\mathbb{R}^{+})=\{0\}.$
Remark 5.
The ultrametric $d^{+}$ on $\mathbb{R}^{+}$ was introduced by Delhommé, Laflamme, Pouzet, and Sauer in [6]. Some results related to the ultrametric space $(\mathbb{R}^{+},d^{+})$ can be found in [20, 12, 22, 21, 23].
4 Center of distances and centered spheres
The next concept was introduced in [13].
Definition 10.
Let $(X,d)$ be a metric space. A set $S\subseteq X$ is a centered sphere in $(X,d)$ if there exist $c\in S$, the center of $S$, and $r\in\mathbb{R}^{+}$, the radius of $S$, such that
$S=\{x\in X:d(x,c)=r\}\cup\{c\}.$ (36)
If $r=0$, then (36) gives the equality $S=\{c\}$. Hence the following proposition holds.
Proposition 5.
Let $(X,d)$ be a metric space. Then the singleton $\{c\}$ is a centered sphere in $(X,d)$ for every $c\in X$.
Theorem 4.
Let $(X,d)\in{\bf UT}$ have at least two points and let $C(X)$ be the center of distances of $(X,d)$. Then the equality
$C(X)=\{0,\operatorname{diam}X\}$ (37)
holds if and only if $X$ is a centered sphere in $(X,d)$ and $\operatorname{diam}X$ is the radius of this sphere.
Proof.
Let $X$ be a centered sphere in $(X,d)$. Then, by Definition 10, there are $c\in X$ and $r\in\mathbb{R}^{+}$ such that
$X=\{x\in X:d(x,c)=r\}\cup\{c\}.$ (38)
The last equality and the strong triangle inequality imply
$d(x,y)\leq\max\{d(x,c),d(c,y)\}\leq r$
for all $x,y\in X$. Consequently we have $\operatorname{diam}X=r$ and $\operatorname{diam}X\in D(X),$ where $D(X)$ is the distance set of $(X,d)$. Hence (37) follows from Theorem 3.
Now we must show that (37) implies (38). Let us do it. Since $(X,d)\in\bf UT$ holds, there is a labeled tree $T=T(l)$ such that $(X,d)$ and $(V(T),d_{l})$ are isometric. Consequently, we have the equalities
$D(X)=D(V(T)),\quad C(X)=C(V(T))$
and can rewrite equality (37) as
$C(V(T))=\{0,\operatorname{diam}V(T)\}.$ (39)
By Lemma 1 we may also suppose that
$D(V(T))=\{l(u)\colon u\in V(T)\}.$ (40)
It follows from (39) that the membership relation
$\operatorname{diam}V(T)\in D(V(T))$
holds, and, consequently, by (40) there exists $u^{*}\in V(T)$ such that
$\operatorname{diam}V(T)=l(u^{*})\geq l(u)$ (41)
for each $u\in V(T)$.
Using formulas (14) and (41), it is easy to prove that
$d_{l}(u^{*},u)=l(u^{*})$
holds for each $u\in V(T)\setminus\{u^{*}\}.$ The last statement implies (38) with $c=\Phi(u^{*})$ and $r=\operatorname{diam}X,$ where $\Phi(u^{*})$ is the image of the point $u^{*}\in V(T)$ under some isometry $\Phi:V(T)\to X$ of the spaces $(V(T),d_{l})$ and $(X,d)$.
The proof is completed. ∎
The next our goal is to describe some conditions under which every open ball of a space $(X,d)\in{\bf UT}$ is a centered sphere in this space.
The following lemma is a reformulation of Lemma 4.3 from [11].
Lemma 3.
Let $(X,d)\in\mathbf{UT}$ and let $T=T(l)$ be a labeled tree such that $(X,d)$ and $(V(T),d_{l})$ are isometric. Then, for every $B^{1}\in{\bf B}_{X}$ there is a subtree $T^{1}$ of $T$ such that $(B^{1},d\big|_{B^{1}\times B^{1}})$ and $(V(T^{1}),d_{l^{1}})$ are isometric, where $l^{1}$ is the restriction of the labeling $l:V(T)\to\mathbb{R}^{+}$ on the set $V(T^{1})$.
We will also use the next lemma.
Lemma 4.
Let $(X,d)$ be an ultrametric space and let $B$ be an open ball in $(X,d)$. Then $B$ is a centered sphere in $(X,d)$ if and only if this ball is a centered sphere in $(B,d|_{B\times B})$.
Proof.
By Definition 10, a ball $B\in{\bf B}_{X}$ is a centered sphere in $(X,d)$ if and only if the equality
$B=\{x\in X\colon d(x,c_{1})=r_{1}\}\cup\{c_{1}\}$ (42)
holds for some $c_{1}\in B$ and $r_{1}\in\mathbb{R}^{+}.$ Similarly, $B$ is a centered sphere in $(B,d|_{B\times B})$ if and only if the equality
$B=\{x\in B:d(x,c_{2})=r_{2}\}\cup\{c_{2}\}$ (43)
holds for some $c_{2}\in B$ and $r_{2}\in\mathbb{R}^{+}$.
Let (42) hold. It follows from $B\subseteq X$ that
$\{x\in X:d(x,c_{1})=r_{1}\}\cup\{c_{1}\}\supseteq\{x\in B:d(x,c_{1})=r_{1}\}\cup\{c_{1}\}.$ (44)
If the last inclusion is strict, then there is $p\in X$ satisfying
$d(p,c_{1})=r_{1}\quad\text{and}\quad p\notin B,$
that contradicts (42). Hence (42) implies the equality.
$\{x\in X:d(x,c_{1})=r_{1}\}\cup\{c_{1}\}=\{x\in B:d(x,c_{1})=r_{1}\}\cup\{c_{1}\},$
that gives us (43) with $r_{2}=r_{1}$ and $c_{2}=c_{1}$. Thus if $B$ is a centered sphere in $(X,d)$, then $B$ also is a centered sphere in $(B,d|_{B\times B})$.
Suppose now that equality (43) holds. We claim that the equality
$\{x\in X:d(x,c_{2})=r_{2}\}\cup\{c_{2}\}=\{x\in B:d(x,c_{2})=r_{2}\}\cup\{c_{2}\}$ (45)
is satisfied.
Similarly (44) we have the inclusion
$\{x\in X:d(x,c_{2})=r_{2}\}\cup\{c_{2}\}\supseteq\{x\in B:d(x,c_{2})=r_{2}\}\cup\{c_{2}\}.$
Consequently, equality (45) holds if we have
$\{x\in X:d(x,c_{2})=r_{2}\}\cup\{c_{2}\}\subseteq\{x\in B:d(x,c_{2})=r_{2}\}\cup\{c_{2}\}.$ (46)
Equality (43) implies $c_{2}\in B$. Hence, by Proposition 1, we may assume
$B=B_{r}(c_{2})=\{x\in X\colon d(x,c_{2})<r\}$ (47)
for some $r>0$. Equalities (43) and (47) imply the inequality $r_{2}<r,$ that together with (47) gives us
$\{x\in X\colon d(x,c_{2})=r_{2}\}\cup\{c_{2}\}\subseteq B.$ (48)
Inclusion (47) follows from (43) and (47).
The proof is completed. ∎
If a subset $S$ of an ultrametric space $(X,d)$ is a centered sphere in itself but $S$ is not a singleton or an open ball in $(X,d)$, then $S$ may not be a centered sphere in $(X,d)$.
$x_{1}$$x_{2}$$x_{3}$$x_{4}$222211222222(X,d)(X,d)22221111P=P(l)P=P(l)
Figure 1: The space $(X,d)$ is isometric to the ultrametric space $(V(P),d_{l})$ generated by labeled path $P=P(l)$.
Example 3.
Let us consider the ultrametric space $(X,d)$ depicted by Figure 1. Then $(X,d)\in{\bf UT}$ and the set $\{x_{1},x_{2}\}$ is a centered sphere in itself but any centered sphere in $(X,d)$ is either a singleton or coincides with one of sets $\{x_{3},x_{4}\}$, $\{x_{1},x_{2},x_{3}\}$, $\{x_{1},x_{2},x_{4}\}$, $\{x_{1},x_{2},x_{3},x_{4}\}$.
In what follows we denote ${\bf Cs}_{X}$ the set of all centered spheres of an ultrametric space $(X,d)$.
Theorem 5.
Let $(X,d)\in{\bf UT}$ and let
$D_{p}(X):=\{d(p,x):x\in X\}$ (49)
for each $p\in X$. Then the following statements are equivalent:
  1. (i)
    The inclusion
    ${\bf B}_{X}\subseteq{\bf Cs}_{X}$ (50)
    holds.
  2. (ii)
    The set $D_{p}(X)\cap[0,r)$ has the greatest element for all $p\in X$ and $r\in(0,\infty)$.
Proof.
$(i)\Rightarrow(ii)$. Let inclusion (50) hold, let $p$ be an arbitrary point of $X$, and let $r\in(0,\infty)$. We must show that the set $D_{p}(X)\cap[0,r)$ has the greatest element. This element obviously exists if $|D_{p}(X)\cap[0,r)|=1$.
Let us now assume that $|D_{p}(X)\cap[0,r)|\geq 2$ hold and consider the open ball
$B:=\{x\in X\colon d(x,p)<r\}.$ (51)
Inclusion (49) implies that $B$ is a centered sphere in $(X,d)$. By Lemma 3 the ultrametric space $(B,d|_{B\times B})$ is an $\bf UT$}-space because $(X,d)\in{\bf UT}$ by condition of the theorem. Now, using Theorem 4 for the space $(B,d|_{B\times B})$, we obtain
$\operatorname{diam}B\in C(B)\subseteq D(B),$
where $C(B)$ and $D(B)$ are the center of distances and, respectively, the distance set of $(B,d|_{B\times B})$. Consequently, the diametrical graph $G_{B}$ of $(B,d|_{B\times B})$ is non-empty. Theorem 1 implies that $G_{B}$ is a complete multipartite graph. Hence, by Definition 5, there is a point $x\in B$ such that
$d(p,x)=\operatorname{diam}B.$
Using (49), (51) and formula (8) with $A=B$, we see that $\operatorname{diam}B$ is the greatest element of the set $D_{p}(X)\cap[0,r)$ as required.
$(ii)\Rightarrow(i)$. Let statement $(ii)$ hold. Let us consider an arbitrary open ball $B^{1}\in{\bf B}_{X}$ with a center $p_{1}\in X$ and a radius $r_{1}>0$,
$B^{1}=\{x\in X\colon d(p_{1},x)<r_{1}\}.$ (52)
To prove inclusion (50) we must show that $B^{1}$ is a centered sphere in $(X,d)$.
If $B^{1}$ is a singleton, then $B^{1}\in{\bf Cs}_{X}$ holds by Proposition 5.
Suppose now that $B^{1}$ contains at least two points. Since $(X,d)$ is an $\bf UT$}-space, Lemma 3 implies the existence of a labeled tree $T^{1}=T^{1}(l_{1})$ such that $(B^{1},d|_{B^{1}\times B^{1}})$ and $(V(T^{1}),d_{l_{1}})$ are isometric. Hence the ultrametric space $(B^{1},d|_{B^{1}\times B^{1}})$ also belongs to the class $\bf UT$}.
By Theorem 4, the membership $(B^{1},d|_{B^{1}\times B^{1}})\in{\bf UT}$ implies $B^{1}\in{\bf Cs}_{B^{1}}$ if and only if
$C(B^{1})=\{0,\operatorname{diam}B^{1}\},$ (53)
where $C(B^{1})$ is the center of distances of $(B^{1},d|_{B^{1}\times B^{1}})$. By Theorem 3, equality (53) is equivalent to
$\operatorname{diam}B^{1}\in D(B^{1}),$ (54)
where $D(B^{1})$ is the distance set of the space $(B^{1},d|_{B^{1}\times B^{1}})$. The space $(B^{1},d|_{B^{1}\times B^{1}})$ is a subspace of $(X,d)$. Hence the inclusion
$D(B^{1})\subseteq D(X)$
holds. Let $D_{p_{1}}$ be the set defined by (49) with $p=p_{1}$, where $p_{1}$ is the center of the ball $B^{1}$. By statement $(ii)$ the set $D_{p_{1}}\cap[0,r_{1})$ has the greatest element $R_{1}$, where $r_{1}$ is the radius of the ball $B^{1}$. Thus we have
$R_{1}\in D_{p_{1}}\cap[0,r_{1})\subseteq D(B^{1}).$
The strong triangle inequality and (52) imply $R_{1}=\operatorname{diam}B^{1}$. Thus (54) holds and, consequently, we have
$B^{1}\in{\bf Cs}_{B^{1}},$ (55)
by Theorems 3 and 4. Using Lemma 4 we may rewrite (55) as
$B^{1}\in{\bf Cs}_{X}.$ (56)
Thus (56) holds for each $B^{1}\in{\bf B}_{X}$. Inclusion (50) follows. The proof is completed. ∎
Corollary 2.
Let $(X,d)\in{\bf UT}$ be totally bounded. Then the inclusion
${\bf B}_{X}\subseteq{\bf Cs}_{X}$
holds.
Proof.
Let $p$ be a point of $X$, let
$D_{p}(X):=\{d(p,x):x\in X\},$
and let $r>0$. The existence of the greatest element in the set $D_{p}(X)\cap[0,r)$ follows from Proposition 2. Hence the inclusion ${\bf B}_{X}\subseteq{\bf Cs}_{X}$ holds by Theorem 5. ∎
Since every compact metric space is totally bounded, Corollary 2 implies the following.
Corollary 3.
Let $(X,d)\in{\bf UT}$ be compact. Then the inclusion
${\bf B}_{X}\subseteq{\bf Cs}_{X}$
holds.
The following example shows that the condition $(X,d)\in{\bf UT}$ cannot be omitted in Theorem 5 in the general case.
Example 4 (The $p$-adic ultrametric).
Let $(X,d)$ be isometric to $({\mathbb{Q}},d_{p})$, where $d_{p}$ is the $p$-adic ultrametric defined in Example 1. Then using formula (7), it is easy to see that statement $(ii)$ of Theorem 5 is valid. Moreover, since the space $({\mathbb{Q}},d_{p})$ does not contain isolated points, there are no open balls in $({\mathbb{Q}},d_{p})$ that are centered spheres in this space. Thus the inclusion ${\bf B}_{X}\subseteq{\bf Cs}_{X}$ is false.
The next theorem is a direct consequence of Theorem 4.9 from [11].
Theorem 6.
Let $(X,d)$ be an ultrametric space. Then the following statements are equivalent:
  1. (i)
    The inclusion ${\bf B}_{X}\supseteq{\bf Cs}_{X}$ holds.
  2. (ii)
    The ultrametric $d$ is equidistant.
  3. (iii)
    The equality ${\bf B}_{X}={\bf Cs}_{X}$ holds.
Example 5.
Let $(X,d)$ be a equidistant metric space. Then each non-empty subset $A$ of $X$ is a centered sphere in $(A,d|_{A\times A})$ but any centered sphere in $(X,d)$ is either a singleton or coincides with the set $X$.
5 Center of distances, diametrical graphs, and weak similarities
Using the concepts of diametrical graphs and star graphs we also can give the necessary and sufficient conditions under which the equality
$C(X)=\{0,\operatorname{diam}X\}$
holds for $\bf UT$}-spaces $(X,d)$.
Theorem 7.
Let $(X,d)\in{\bf UT}$ have at least two points and let $G_{X}$ be the diametrical graph of $(X,d)$. Then the following statements are equivalent:
  1. (i)
    The equality
    $C(X)=\{0,\operatorname{diam}X\}$ (57)
    holds.
  2. (ii)
    $G_{X}$ is a complete multipartite graph such that at least one of its parts is a singleton.
  3. (iii)
    $G_{X}$ has a spanning star subgraph.
Proof.
$(i)\Rightarrow(ii)$. Let (57) hold. Then the membership relation
$\operatorname{diam}X\in D(X)$ (58)
is valid, because $C(X)$ is a subset of $D(X)$.
Definition 6 and membership relation (58) imply that the graph $G_{X}$ is non-empty. Consequently, this graph is complete multipartite by Theorem 1. By Theorem 4, equality (57) holds if and only if $X$ is a centered sphere in $(X,d)$ and $\operatorname{diam}X$ is the radius of this sphere.
Let $c$ be the center of the centered sphere $X$. Then, using (36) with $S=X$ and $r=\operatorname{diam}X$ we obtain the equality
$d(x,c)=\operatorname{diam}X$ (59)
for each $x\in X\setminus\{c\}$. Hence, by Definition 6, we have
$\{x,c\}\in E(G_{X})$ (60)
for each $x\in X\setminus\{c\}$. Thus, the singleton $\{c\}$ is a part of $G_{X}$ by Definition 5. Consequently, the implication $(i)\Rightarrow(ii)$ is valid.
$(ii)\Rightarrow(i)$. Let $(ii)$ hold. Then $G_{X}$ is a complete multipartite graph and, consequently, $G_{X}$ is non-empty by Definition 5. Hence we can find two distinct points $x,y\in X$ such that
$\{x,y\}\in E(G_{X}).$ (61)
By Definition 6, membership (61) implies the equality
$d(x,y)=\operatorname{diam}X.$
Thus we have
$\operatorname{diam}X\in D(X),$
that implies statement $(i)$ by Theorem 3.
$(ii)\Rightarrow(iii)$. Let a singleton $\{c\}$ be a part of the complete multipartite graph $G_{X}$. Then the membership relation
$\{x,c\}\in E(G_{X})$
holds for all $x\in X\setminus\{c\}$ by Definition 5. Let $St_{X}$ be a subgraph of $G_{X}$ such that
$V(St_{X}):=X$ (62)
and
$E(St_{X}):=\{\{x,c\}\colon x\in X\setminus\{c\}\}.$ (63)
Then $St_{X}$ is a star graph by Definition 8 and, moreover, $St_{X}$ is a spanning subtree of $G_{X}$ by (62). Thus $(ii)\Rightarrow(iii)$ is a valid implication.
$(iii)\Rightarrow(ii)$. Let $St_{X}$ be a spanning star subgraph of the diametrical graph $G_{X}$. Then (62) holds. Since $(X,d)$ contains at least two points, $St_{X}$ is non-empty by Definition 8. Thus $G_{X}$ also is non-empty by inclusion $St_{X}\subseteq G_{X}$. Consequently, $G_{X}$ is complete multipartite by Theorem 1. Definition 8 and equality (62) imply that there is $c\in X$ such that (63) holds for each $x\in X\setminus\{c\}.$ Consequently the singleton $\{c\}$ is a part of the complete multipartite graph $G_{X}$. Thus $(iii)$ implies $(ii)$.
The proof is completed. ∎
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$St_{X}$$x_{1}$$x_{2}$$x_{3}$$x_{4}$$G_{X}$
Figure 2: A spanning star graph $St_{X}$ of the diametrical graph $G_{X}$ of the space $(X,d)$ depicted by Figure 1.
Remark 6.
Analyzing the proof of Theorem 7 we can see that the equivalence $(ii)\Leftrightarrow(iii)$ is valid for arbitrary complete multipartite graph $G$. Consequently, Theorem 7 gives us necessary and sufficient conditions under which complete multipartite graphs have spanning star subgraphs. In connection with this result, we note that all connected graphs have spanning subtrees. See, for example, Proposition 14 in Chapter 1 of book [28].
The next concept was introduced in [14] as a generalization of the concept of similarity for general semimetric spaces.
Definition 11.
Let $(X,d)$, $(Y,\rho)$ be metric spaces and let $D(X)$, $D(Y)$ be the distance sets of $(X,d)$, $(Y,\rho)$ respectively. A bijective mapping $\Phi\colon X\to Y$ is a weak similarity if there is a strictly increasing bijection $f\colon D(Y)\to D(X)$ such that
$d(x,y)=f\left(\rho\left(\Phi(x),\Phi(y)\right)\right)$
for all $x,y\in X$.
We will say that metric spaces $(X,d)$ and $(Y,\rho)$ are weakly similar if there is a weak similarity $X\to Y$.
The next theorem is a reformulation of Theorem 3.22 from [5].
Theorem 8.
Let $(X,d)$ be an ultrametric space with $|X|\geq 2.$ Then the following statements are equivalent:
  1. (i)
    $(X,d)$ is weakly similar to an unbounded ultrametric space.
  2. (ii)
    The distance set $D(X)$ does not contain $\operatorname{diam}X$,
    $\operatorname{diam}X\notin D(X).$
The last theorem and Theorem 3 give us the following corollary.
Corollary 4.
Let $(X,d)\in{\bf UT}$ have at least two points and let $C(X)$ be the center of distances of $(X,d)$. Then the equality
$C(X)=\{0\}$ (64)
holds if and only if $(X,d)$ is weakly similar to an unbounded ultrametric space.
Proof.
Let $D(X)$ be the distance set of $(X,d)$. Then, by Theorem 3, equality (64) holds iff
$\operatorname{diam}X\notin D(X).$
It implies the desired result by Theorem 8. ∎
Remark 7.
Example 1 shows that equality (64) may not hold if $(X,d)\notin{\bf UT}$.
6 Conclusion. Expected results
Let $(X,d)$ be a metric space. A closed ball in $(X,d)$ is a subset $\overline{B}_{r}(c)$ of the set $X$ defined as
$\overline{B}_{r}(c):=\{x\in X\colon d(c,x)\leq r\},$
where $c\in X$ and $r\in\mathbb{R}^{+}$.
Let us denote by $\overline{\bf B}_{X}$ the set of all closed balls of $(X,d)$.
It follows directly from Definition 10, that every centered sphere in the space $(X,d)$ is a closed ball of this space,
${\bf Cs}_{X}\subseteq\overline{\bf B}_{X}.$
The following conjecture is a counterpart of Theorem 5.
Conjecture 1.
Let $(X,d)\in{\bf UT}$ and let
$D_{p}(X):=\{d(p,x):x\in X\}$
for each $p\in X$. Then the following statements are equivalent:
  1. (i)
    The inclusion
    $\overline{{\bf B}}_{X}\subseteq{\bf Cs}_{X}$
    holds.
  2. (ii)
    The set $D_{p}(X)\cap[0,r]$ has the greatest element for all $p\in X$ and $r\in\mathbb{R}^{+}$.
Let $(X_{3},d)$ be the three-point ultrametric space depicted in Figure 3 below.
[…]
Figure 3: A non-equidistant three-point ultrametric space $(X_{3},d)$.
Then direct calculations show that every non-empty subset of $X_{3}$ is a centered sphere in $(X_{3},d)$.
Conjecture 2.
Let $(Y,\rho)$ be an ultrametric space such that $|Y|\geq 3$. Then the following statements are equivalent:
  1. (i)
    Each non-empty subset of the set $Y$ is a centered sphere in $(Y,\rho)$.
  2. (ii)
    The set $Y$ contains exactly three points and $(Y,\rho)$ is weakly similar to $(X_{3},d)$.
The following conjecture was motivated by properties of $p$-adic ultrametric with $p=2$ (see Example 1).
Conjecture 3.
Let $n\geq 1$ be an integer number, let $\log_{2}(n)$ be the binary logarithm of $n$, and let $\lfloor\log_{2}(n)\rfloor$ be the integer part of $\log_{2}(n)$. Then the inequality
$C(X)\leq 1+\lfloor\log_{2}(n)\rfloor$
holds for each ultrametric space $(X,d)$ with $|X|=n$. Moreover, there exists an ultrametric space $(Y,\rho)$ such that $|Y|=n$ and
$C(Y)=1+\lfloor\log_{2}(n)\rfloor.$
Acknowledgement
The first author was supported by grant 367319 of the Research Council of Finland.

### v2

@@
 The center of distances of a metric space $(X,d)$ is the set $C(X)$ of all $t\in\mathbb{R}^{+}$ for which the equation $d(x,p)=t$ has a solution for each $p\in X$.
-We prove that the equalities $C(X)=\{0\}$ or $C(X)=\{\operatorname{diam}X,0\}$ hold if $(X,d)$ is an ultrametric space generated by labeled trees.
+We prove that the equalities $C(X)=\{0\}$ or $C(X)=\{0,\operatorname{diam}X\}$ hold if $(X,d)$ is an ultrametric space generated by labeled trees.
 The necessary and sufficient conditions under which $\operatorname{diam}X\in C(X)$ are found.
@@
 It was shown in [8] that the compact ultrametric spaces generated by labeled star graphs are the completions of totally bounded ultrametric space generated by decreasingly labeled rays.
-The labeled trees generating separable ultrametric spaces are characterized, up to isomorphism, in [15].
 In the present paper we investigate the class $\bf UT$} of all ultrametric spaces which are isometric to spaces generated by labeled trees.
-The main goal of the paper is the characterization of the center of distances $C(X)$ of $\bf UT$}-spaces $(X,d)$.
+The main goal of the paper is the characterization of the center of distances of $\bf UT$}-spaces $(X,d)$.
+The notion of center of distances was introduced in paper [2] and used for a generalization of John von Neumann Theorem on permutations of two sequences with the same set of cluster points [22].
+At the moment the authors are not aware of any results describing the structure of the center of distances in the case of ultrametric spaces.
 The paper is organized as follows. The next section contains some definitions and facts from the theory of ultrametric spaces and graph theory. The main results of the paper are proved in Sections 35.
-Theorem 3 of Section 3 shows that the center of distances $C(X)$ of arbitrary $\bf UT$}-space $(X,d)$ satisfies either the equality
+Theorem 4 of Section 3 shows that the center of distances $C(X)$ of an arbitrary $\bf UT$}-space $(X,d)$ satisfies either the equality
 $C(X)=\{0,\operatorname{diam}X\}$ (1)
@@
 Theorem 8 of Section 5 shows that equality (1) holds for $\bf UT$}-space $(X,d)$ if and only if the diametrical graph of $(X,d)$ has a spanning star subgraph.
-Corollary 4 of Theorem 8 of Section 5 shows that equality (2) holds for $\bf UT$}-space $(X,d)$ with $|X|\geq 2$ if and only if $(X,d)$ is weakly similar to an unbounded ultrametric space.
+Corollary 4 of Theorem 1 shows that equality (2) holds for $\bf UT$}-space $(X,d)$ with $|X|\geq 2$ if and only if $(X,d)$ is weakly similar to an unbounded ultrametric space.
+The final Section 6 contains some conjectures related to results of the paper.
 2 Preliminaries
@@
 holds for all $x$, $y$, $z\in X$. The function $d$ is called an ultrametric on $X$.
-Following [7, p. 46] we say that a metric $d:X\times X\to\mathbb{R}^{+}$ is equidistant if there is a constant $k>0$ such that
+Following [5, p. 46] we say that a metric $d:X\times X\to\mathbb{R}^{+}$ is equidistant if there is a constant $k>0$ such that
 $d(x,y)=k$
@@
 $D(X):=\{d(x,y):x,y\in X\}.$ (3)
-The concept of center of distances was introduced in [4] as follows.
 Definition 3.
@@
 $C(X):=\{t\in D(X):\forall p\in X\ \exists x\in X\ d(p,x)=t\}.$ (4)
-Remark 1.
-Some properties of the center of distances are described in [2, 3, 24, 25].
 Example 1 (The $p$-adic ultrametric).
@@
 is the diameter of $A$. If the inequality $\operatorname{diam}A<\infty$ holds, then we say that $A$ is a bounded subset of $(X,d)$.
+The next concept was introduced in [11] as a generalization of the concept of similarity for general semimetric spaces.
+Definition 4.
+Let $(X,d)$, $(Y,\rho)$ be metric spaces and let $D(X)$, $D(Y)$ be the distance sets of $(X,d)$, $(Y,\rho)$ respectively.
+A bijective mapping $\Phi\colon X\to Y$ is a weak similarity if there is a strictly increasing bijection $f\colon D(Y)\to D(X)$ such that
+$d(x,y)=f\left(\rho\left(\Phi(x),\Phi(y)\right)\right)$
+for all $x,y\in X$.
+We will say that metric spaces $(X,d)$ and $(Y,\rho)$ are weakly similar if there is a weak similarity $X\to Y$.
+The next theorem is a reformulation of Theorem 3.22 from [3].
+Theorem 1.
+Let $(X,d)$ be an ultrametric space with $|X|\geq 2.$ Then the following statements are equivalent:
+  1. (i)
+    $(X,d)$ is weakly similar to an unbounded ultrametric space.
+  2. (ii)
+    The distance set $D(X)$ does not contain $\operatorname{diam}X$,
+    $\operatorname{diam}X\notin D(X).$
+Lemma 1.
+Let $A$ be a non-empty subset of an ultrametric space $(X,d)$. Then the equality
+$\operatorname{diam}A=\sup\{d(p,a):a\in A\}$ (9)
+holds for each $p\in A$.
+Proof.
+Let $p$ be an arbitrary point of $A$. Then, using the strong triangle inequality, we obtain
+$d(x,y)\leq\max\{d(p,x),d(p,y)\}\leq\sup\{d(p,a)\colon a\in A\}.$
+Consequently, the inequality
+$\operatorname{diam}A\leq\sup\{d(p,a)\colon a\in A\}$
+holds. The converse inequality follows directly from (8). Thus equality (9) holds. ∎
+Remark 1.
+Equality (9) is known, see, for example, equality (1.3) in [15], but the authors cannot give here an exact reference to standard books on ultrametric spaces.
 Let $(X,d)$ be a metric space. The open ball with a radius $r>0$ and a center $c\in X$ is the set
@@
 A graph $G$ is empty if $E(G)=\emptyset$.
-The following concept is well-known for finite graphs (see, for example, [8, p. 17]).
+The following concept is well-known for finite graphs (see, for example, [6, p. 17]).
 Definition 6.
@@
 We say that $G$ is a complete multipartite graph if $G$ is complete $k$-partite for some $k$.
-Our next definition is a modification of Definition 2.1 from [26].
+The next concept was introduced in [19] for characterization of finite ultrametric spaces for which the Gomory–Hu inequality [17] turns to equality.
 Definition 7.
@@
     The diametrical graph $G_{X}$ is complete multipartite.
-Corollary 1.
-Let $(X,d)$ be an ultrametric space with $|X|\geq 2$, let $D(X)$ be the distance set of $(X,d)$, and let $C(X)$ be the center of distances of $(X,d)$. If $\operatorname{diam}X$ belongs to $D(X)$,
-$\operatorname{diam}X\in D(X),$ (11)
-then $\operatorname{diam}X$ also belongs to $C(X)$,
-$\operatorname{diam}X\in C(X).$ (12)
-Proof.
-If (11) holds, then the diametrical graph $G_{X}$ is non-empty by Remark 2. Hence $G_{X}$ is complete multipartite by Theorem 1. Now Definitions 3 and 5 imply (12). ∎
 Let $G$ be a graph. A graph $G_{1}$ is a subgraph of $G$ if
@@
 with $k\geqslant 2$.
-We will write $P=(x_{1},\ldots,x_{k})$ or $P=P_{x_{1},x_{k}}$ if $P$ is a path satisfying (13) and said that $P$ is a path joining $x_{1}$ and $x_{k}$.
+We will write $P=(x_{1},\ldots,x_{k})$ or $P=P_{x_{1},x_{k}}$ if $P$ is a path satisfying (12) and say that $P$ is a path joining $x_{1}$ and $x_{k}$.
 A graph $G$ is connected if for every two distinct vertices of $G$ there is a path $P\subseteq G$ joining these vertices.
@@
 $\max\{l(u^{\circ}),l(v^{\circ})\}=\max\{l^{*}(u^{\circ}),l^{*}(v^{\circ})\}.$ (20)
-The last equality and formula (18) show that $l^{*}\colon V(T)\to\mathbb{R}^{+}$ is a non-degenerate labeling of $V(T)$ because $l\colon V(T)\to\mathbb{R}^{+}$ is non-degenarate.
+The last equality and formula (17) show that $l^{*}\colon V(T)\to\mathbb{R}^{+}$ is a non-degenerate labeling of $V(T)$ because $l\colon V(T)\to\mathbb{R}^{+}$ is non-degenerate.
 Thus $(V(T),d_{l^{*}})$ is an ultrametric space by Theorem 3.
@@
 Propositions 3 and 4 are known, see, for example, [18] and [2].
+Lemma 1 and Proposition 3 imply the following corollary.
+Corollary 1.
+Let $(X,d)$ be an ultrametric space, $D(X)$ be the distance set of $(X,d)$ and $C(X)$ be the centers of distances of $(X,d)$. If $\operatorname{diam}X$ belongs to $D(X)$,
+$\operatorname{diam}X\in D(X),$
+then $\operatorname{diam}X$ also belongs to $C(X)$,
+$\operatorname{diam}X\in C(X).$
 Lemma 3.
@@
 The ultrametric $d^{+}$ on $\mathbb{R}^{+}$ was introduced by Delhommé, Laflamme, Pouzet, and Sauer in [4].
-Some results related to the ultrametric space $(\mathbb{R}^{+},d^{+})$ can be found in [20, 12, 22, 21, 23].
 4 Center of distances and centered spheres
@@
 The proof is completed. ∎
-The next our goal is to describe some conditions under which every open ball of a space $(X,d)\in{\bf UT}$ is a centered sphere in this space.
+Our next goal is to describe some conditions under which every open ball of a space $(X,d)\in{\bf UT}$ is a centered sphere in this space.
 The following lemma is a reformulation of Lemma 4.3 from [9].
@@
 is satisfied.
-Similarly (44) we have the inclusion
+Similar to (43), we have the inclusion
 $\{x\in X:d(x,c_{2})=r_{2}\}\cup\{c_{2}\}\supseteq\{x\in B:d(x,c_{2})=r_{2}\}\cup\{c_{2}\}.$
@@
 Let us consider the ultrametric space $(X,d)$ depicted by Figure 1.
-Then $(X,d)\in{\bf UT}$ and the set $\{x_{1},x_{2}\}$ is a centered sphere in itself but any centered sphere in $(X,d)$ is either a singleton or coincides with one of sets $\{x_{3},x_{4}\}$, $\{x_{1},x_{2},x_{3}\}$, $\{x_{1},x_{2},x_{4}\}$, $\{x_{1},x_{2},x_{3},x_{4}\}$.
+Then $(X,d)\in{\bf UT}$ and the set $\{x_{1},x_{2}\}$ is a centered sphere in itself but this set is not a centered sphere in $(X,d)$.
+Indeed, the equality $d(x_{1},x_{2})=2$ and the definition of $(X,d)$ give us the equalities
+$\{x\in X\colon d(x,x_{1})=2\}\cup\{x_{1}\}=\{x_{1},x_{2},x_{3},x_{4}\}=\{x\in X\colon d(x,x_{2})=2\}\cup\{x_{2}\}.$
 In what follows we denote ${\bf Cs}_{X}$ the set of all centered spheres of an ultrametric space $(X,d)$.
 Theorem 6.
-Let $(X,d)\in{\bf UT}$ and let
+Let $(X,d)$ be an ultrametric space and let
 $D_{p}(X):=\{d(p,x):x\in X\}$ (48)
-for each $p\in X$. Then the following statements are equivalent:
-  1. (i)
-    The inclusion
+for each $p\in X$. If the inclusion
 ${\bf B}_{X}\subseteq{\bf Cs}_{X}$ (49)
-    holds.
-  2. (ii)
-    The set $D_{p}(X)\cap[0,r)$ has the greatest element for all $p\in X$ and $r\in(0,\infty)$.
+holds, then the set $D_{p}(X)\cap[0,r)$ has the greatest element for all $p\in X$ and $r\in(0,\infty)$.
+Moreover, if $(X,d)\in{\bf UT}$ and the set $D_{p}(X)\cap[0,r)$ has the greatest element for all $p\in X$ and $r\in(0,\infty)$, then inclusion (49) holds.
 Proof.
-$(i)\Rightarrow(ii)$.
-Let inclusion (50) hold, let $p$ be an arbitrary point of $X$, and let $r\in(0,\infty)$.
-We must show that the set $D_{p}(X)\cap[0,r)$ has the greatest element.
-This element obviously exists if $|D_{p}(X)\cap[0,r)|=1$.
-Let us now assume that $|D_{p}(X)\cap[0,r)|\geq 2$ hold and consider the open ball
+Let inclusion (49) hold, let $p$ be an arbitrary point of $X$, and let $r\in(0,\infty)$. We must show that the set $D_{p}(X)\cap[0,r)$ has the greatest element.
+Let us consider the open ball
 $B:=\{x\in X\colon d(x,p)<r\}.$ (50)
-Inclusion (49) implies that $B$ is a centered sphere in $(X,d)$.
-By Lemma 3 the ultrametric space $(B,d|_{B\times B})$ is an $\bf UT$}-space because $(X,d)\in{\bf UT}$ by condition of the theorem.
-Now, using Theorem 4 for the space $(B,d|_{B\times B})$, we obtain
-$\operatorname{diam}B\in C(B)\subseteq D(B),$
-where $C(B)$ and $D(B)$ are the center of distances and, respectively, the distance set of $(B,d|_{B\times B})$.
+By inclusion (49) the ball $B$ is a centered sphere in $(X,d)$, i.e., the equality
+$B=\{x\in X\colon\ d(x,p_{1})=r_{1}\}\cup\{p_{1}\}$
+holds with some $p_{1}\in B$ and $r_{1}\in\mathbb{R}^{+}$. The last equality and the Lemma 1 with $A=B$ and $p=p_{1}$ imply
+$\operatorname{diam}B=r_{1}\quad\text{and}\quad\operatorname{diam}B\in D(B),$
+where $D(B)$ is the distance set of $(B,d|_{B\times B})$.
 Consequently, the diametrical graph $G_{B}$ of $(B,d|_{B\times B})$ is non-empty.
@@
 Using (48), (50) and formula (8) with $A=B$, we see that $\operatorname{diam}B$ is the greatest element of the set $D_{p}(X)\cap[0,r)$ as required.
-$(ii)\Rightarrow(i)$. Let statement $(ii)$ hold. Let us consider an arbitrary open ball $B^{1}\in{\bf B}_{X}$ with a center $p_{1}\in X$ and a radius $r_{1}>0$,
+Let $(X,d)\in{\bf UT}$ and let the set $D_{p}(X)\cap[0,r)$ contain the greatest element for all $p\in X$ and $r\in(0,\infty)$.
+Let us consider an arbitrary open ball $B^{1}\in{\bf B}_{X}$ with a center $p_{1}\in X$ and a radius $r_{1}>0$,
 $B^{1}=\{x\in X\colon d(p_{1},x)<r_{1}\}.$ (51)
@@
 holds.
-Let $D_{p_{1}}$ be the set defined by (49) with $p=p_{1}$, where $p_{1}$ is the center of the ball $B^{1}$.
-By statement $(ii)$ the set $D_{p_{1}}\cap[0,r_{1})$ has the greatest element $R_{1}$, where $r_{1}$ is the radius of the ball $B^{1}$.
+Let $D_{p_{1}}(X)$ be the set defined by (48) with $p=p_{1}$, where $p_{1}$ is the center of the ball $B^{1}$.
+By supposition the set $D_{p_{1}}(X)\cap[0,r_{1})$ has the greatest element $R_{1}$.
 Thus we have
-$R_{1}\in D_{p_{1}}\cap[0,r_{1})\subseteq D(B^{1}).$
-The strong triangle inequality and (52) imply $R_{1}=\operatorname{diam}B^{1}$. Thus (54) holds and, consequently, we have
+$R_{1}\in D_{p_{1}}(X)\cap[0,r_{1})\subseteq D(B^{1}).$
+Lemma 1 and (51) imply $R_{1}=\operatorname{diam}B^{1}$. Thus (53) holds and, consequently, we have
 $B^{1}\in{\bf Cs}_{B^{1}},$ (54)
-by Theorems 3 and 4. Using Lemma 4 we may rewrite (55) as
+by Theorems 4 and 5. Using Lemma 5 we may rewrite (54) as
 $B^{1}\in{\bf Cs}_{X}.$ (55)
-Thus (56) holds for each $B^{1}\in{\bf B}_{X}$. Inclusion (50) follows. The proof is completed. ∎
+Thus (55) holds for each $B^{1}\in{\bf B}_{X}$. Inclusion (49) follows.
+The proof is completed. ∎
 Corollary 2.
@@
 Example 4 (The $p$-adic ultrametric).
-Let $(X,d)$ be isometric to $({\mathbb{Q}},d_{p})$, where $d_{p}$ is the $p$-adic ultrametric defined in Example 1.
-Then using formula (7), it is easy to see that statement $(ii)$ of Theorem 5 is valid.
-Moreover, since the space $({\mathbb{Q}},d_{p})$ does not contain isolated points, there are no open balls in $({\mathbb{Q}},d_{p})$ that are centered spheres in this space.
+Let $(X,d)$ be isometric to $({\mathbb{Q}},d_{p})$, where $d_{p}$ is the $p$-adic ultrametric defined in Example 1 and let
+$D_{q}(X):=\{d(q,x)\colon x\in X\}$
+for $q\in X$.
+Then using formula (7), it is easy to see that the set $D_{q}(X)\cap[0,r)$ contains the greatest element for all $q\in X$ and $r\in(0,\infty)$.
+Since the space $({\mathbb{Q}},d_{p})$ does not contain isolated points, there are no open balls in $({\mathbb{Q}},d_{p})$ that are centered spheres in this space.
 Thus the inclusion ${\bf B}_{X}\subseteq{\bf Cs}_{X}$ is false.
-The next theorem is a direct consequence of Theorem 4.9 from [11].
+It should be noted that singletons are centered spheres by Proposition 5 but such sets cannot be open balls if the space does not contain isolated points.
+It is interesting to compare Theorem 6 with the following result.
 Theorem 7.
@@
     The equality ${\bf B}_{X}={\bf Cs}_{X}$ holds.
+For the proof of Theorem 7 see Theorem 4.9 from [9].
 Example 5.
-Let $(X,d)$ be a equidistant metric space.
+Let $(X,d)$ be an equidistant metric space.
 Then each non-empty subset $A$ of $X$ is a centered sphere in $(A,d|_{A\times A})$ but any centered sphere in $(X,d)$ is either a singleton or coincides with the set $X$.
@@
 Let $c$ be the center of the centered sphere $X$. Then, using (35) with $S=X$ and $r=\operatorname{diam}X$ we obtain the equality
-$d(x,c)=\operatorname{diam}X$ (59)
+$d(x,c)=\operatorname{diam}X$
 for each $x\in X\setminus\{c\}$. Hence, by Definition 7, we have
-$\{x,c\}\in E(G_{X})$ (60)
+$\{x,c\}\in E(G_{X})$
 for each $x\in X\setminus\{c\}$. Thus, the singleton $\{c\}$ is a part of $G_{X}$ by Definition 6. Consequently, the implication $(i)\Rightarrow(ii)$ is valid.
@@
 See, for example, Proposition 14 in Chapter 1 of book [21].
-The next concept was introduced in [14] as a generalization of the concept of similarity for general semimetric spaces.
-Definition 11.
-Let $(X,d)$, $(Y,\rho)$ be metric spaces and let $D(X)$, $D(Y)$ be the distance sets of $(X,d)$, $(Y,\rho)$ respectively.
-A bijective mapping $\Phi\colon X\to Y$ is a weak similarity if there is a strictly increasing bijection $f\colon D(Y)\to D(X)$ such that
-$d(x,y)=f\left(\rho\left(\Phi(x),\Phi(y)\right)\right)$
-for all $x,y\in X$.
-We will say that metric spaces $(X,d)$ and $(Y,\rho)$ are weakly similar if there is a weak similarity $X\to Y$.
-The next theorem is a reformulation of Theorem 3.22 from [5].
-Theorem 8.
-Let $(X,d)$ be an ultrametric space with $|X|\geq 2.$ Then the following statements are equivalent:
-  1. (i)
-    $(X,d)$ is weakly similar to an unbounded ultrametric space.
-  2. (ii)
-    The distance set $D(X)$ does not contain $\operatorname{diam}X$,
-    $\operatorname{diam}X\notin D(X).$
-The last theorem and Theorem 3 give us the following corollary.
+Theorems 1 and 4 give us the following.
 Corollary 4.
@@
 Let us denote by $\overline{\bf B}_{X}$ the set of all closed balls of $(X,d)$.
-It follows directly from Definition 10, that every centered sphere in the space $(X,d)$ is a closed ball of this space,
-${\bf Cs}_{X}\subseteq\overline{\bf B}_{X}.$
 The following conjecture is a counterpart of Theorem 6.
 Conjecture 1.
-Let $(X,d)\in{\bf UT}$ and let
+Let $(X,d)$ be an ultrametric space and let
 $D_{p}(X):=\{d(p,x):x\in X\}$
-for each $p\in X$. Then the following statements are equivalent:
-  1. (i)
-    The inclusion
-    $\overline{{\bf B}}_{X}\subseteq{\bf Cs}_{X}$
-    holds.
-  2. (ii)
-    The set $D_{p}(X)\cap[0,r]$ has the greatest element for all $p\in X$ and $r\in\mathbb{R}^{+}$.
+for each $p\in X$. If the inclusion
+$\overline{{\bf B}}_{X}\subseteq{\bf Cs}_{X}$ (62)
+holds, then the set $D_{p}(X)\cap[0,r]$ has the greatest element for all $p\in X$ and $r\in\mathbb{R}^{+}$.
+Moreover, if $(X,d)\in{\bf UT}$ and the set $D_{p}(X)\cap[0,r]$ has the greatest element for all $p\in X$ and $r\in\mathbb{R}^{+}$, then inclusion (62) holds.
 Let $(X_{3},d)$ be the three-point ultrametric space depicted in Figure 3 below.
@@
 Figure 3: A non-equidistant three-point ultrametric space $(X_{3},d)$.
-Then direct calculations show that every non-empty subset of $X_{3}$ is a centered sphere in $(X_{3},d)$.
+Then it is easy to show that every non-empty subset of $X_{3}$ is a centered sphere in $(X_{3},d)$.
 Conjecture 2.
@@
     The set $Y$ contains exactly three points and $(Y,\rho)$ is weakly similar to $(X_{3},d)$.
-The following conjecture was motivated by properties of $p$-adic ultrametric with $p=2$ (see Example 1).
+The following conjecture was motivated by some calculations with $p$-adic ultrametric with $p=2$ (see Example 1).
 Conjecture 3.
 Let $n\geq 1$ be an integer number, let $\log_{2}(n)$ be the binary logarithm of $n$, and let $\lfloor\log_{2}(n)\rfloor$ be the integer part of $\log_{2}(n)$. Then the inequality
-$C(X)\leq 1+\lfloor\log_{2}(n)\rfloor$
+$|C(X)|\leq 1+\lfloor\log_{2}(n)\rfloor$
 holds for each ultrametric space $(X,d)$ with $|X|=n$. Moreover, there exists an ultrametric space $(Y,\rho)$ such that $|Y|=n$ and
-$C(Y)=1+\lfloor\log_{2}(n)\rfloor.$
-Acknowledgement
+$|C(Y)|=1+\lfloor\log_{2}(n)\rfloor.$
+To see why a binary logarithm bound is expected in Conjecture 3, we only note that the following simple statement is valid.
+Let $\{X_{1},X_{2},\dots,X_{k}\}$ be a partition of a finite set $X$ on disjoint subsets $X_{1},X_{2},\dots,X_{k}$ with $k\geq 2$. Then there is a partition $\{X_{1}^{*},X_{2}^{*}\}$ of $X$ such that
+$\min\{|X_{1}^{*}|,|X_{2}^{*}|\}\geq\min\{|X_{1}|,\dots,|X_{k}|\}.$
+Funding
 The first author was supported by grant 367319 of the Research Council of Finland.